# Concurrent Composition for Differentially Private Continual Mechanisms

### v3

@@ Acknowledgments
1Monika Henzinger and Roodabeh Safavi were supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101019564), and the Austrian Science Fund (FWF) under grants DOI 10.55776/Z422, DOI 10.55776/I5982, and DOI 10.55776/P33775, with additional funding from the netidee SCIENCE Stiftung, 2020–2024.
2Salil Vadhan was supported by NSF grant BCS-2218803, a grant from the Sloan Foundation, and a Simons Investigator Award. Work began while a Visiting Researcher at the Bocconi University Department of Computing Sciences, supported by Luca Trevisan’s ERC Project GA-834861.

### v4

@@ Acknowledgments
1Salil Vadhan was supported by NSF grant BCS-2218803, a grant from the Sloan Foundation, and a Simons Investigator Award. Work began while a Visiting Researcher at the Bocconi University Department of Computing Sciences, supported by Luca Trevisan’s ERC Project GA-834861.
2Monika Henzinger and Roodabeh Safavi were supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101019564), and the Austrian Science Fund (FWF) under grants DOI 10.55776/Z422, DOI 10.55776/I5982, and DOI 10.55776/P33775. For open access purposes, the author has applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission.
Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Figures
[…]
Comparison between interactive mechanism (IMs), continual observation (CO), and continual mechanisms (CMs)
Appendix B Variants of Concurrent Composition
Here, we elaborate on the different types of concurrent composition that we study in this paper: